determine the milliliters of 10% sulfuric acid and the milliliters of 25% sulfuric acid that should be mixed to obtain 20 milliliters of 18% sulfuric acid
Answer by richwmiller(17219)

x+y=20,
0.1*x+0.25*y=0.18*20
0.1*x+0.25*y=3.6
x=20-y
0.1*(20-y)+0.25*y=3.6
2-0.1y+0.25*y=3.6
0.15*y=1.6
y=10.6666667 at 25%
x=20-y
x=9.33333333 at 10%
check
0.1*9.33333333+0.25*10.6666667=0.18*20
0.93333333+2.66666667=3.6
3.6=3.6
ok